# Progress-Aware Video Frame Captioning

### v1

@@ 2.1 Experimental setup
For the user study, 15 participants reviewed 85 frame sequences (364 frames) in total, randomly sampled from the FrameEvalBenchmark. We evaluate the captions from four leading models—two open-source (LLAVA-OV [33], Qwen2-VL [62]) and two proprietary (Gemini-1.5-Pro [53], GPT-4o [2])—alongside our ProgressCaptioner. Note that image captioning baselines are excluded due to their excessively lengthy captions and complete lack of temporal coherence. Participants are presented with captions produced by these five models, randomly shuffled for each sequence, and asked to choose the best and second best (with an additional “none” option available) for each frame’s caption. The average selection rate per model is reported, providing insights into subjective caption quality preferences.
@@
For results in Section 4.1, direct inference is applied on $T$ frames. For results in Section 4.2, we employ a 2-frame sliding window, where ProgressCaptioner performs frame pair captioning (except for NeXT-QA, where we uniformly sample 6 frames from the original video and apply direct inference on these 6-frame sequences without a sliding window). A single frame ($v_{t}$) can receive two captions: one from the pair ($v_{t-1}$, $v_{t}$) and another from ($v_{t}$, $v_{t+1}$). We concatenate the two captions for frame classification tasks to provide richer contextual information, aiding the LLM in frame label prediction. For keyframe selection, we use the caption from the pair ($v_{t-1}$, $v_{t}$) for $v_{t}$ to maintain caption sequence coherence.

### v2

@@ 2.1 Experimental setup
User Study
To evaluate the subjective quality of generated captions, we conduct a user study involving 15 graduate student participants fluent in English. The study utilized 85 randomly sampled frame sequences (totaling 364 frames) from the FrameCapEval benchmark. We evaluate the captions from four leading models—two open-source (LLAVA-OV [33], Qwen2-VL [62]) and two proprietary (Gemini-1.5-Pro [53], GPT-4o [2])—alongside our ProgressCaptioner. Note that image captioning baselines are excluded due to their excessively lengthy captions and complete lack of temporal coherence. Participants are presented with captions produced by these five models, randomly shuffled for each sequence, and asked to choose the best and second best (with an additional “none” option available) for each frame’s caption. The average selection rate per model is reported, providing insights into subjective caption quality preferences.
@@
For results in Section 4.1, direct inference is applied on $T$ frames. For results in Section LABEL:sec:exp_application, we employ a 2-frame sliding window, where ProgressCaptioner performs frame pair captioning (except for NeXT-QA, where we uniformly sample 6 frames from the original video and apply direct inference on these 6-frame sequences without a sliding window). A single frame ($v_{t}$) can receive two captions: one from the pair ($v_{t-1}$, $v_{t}$) and another from ($v_{t}$, $v_{t+1}$). We concatenate the two captions for frame classification tasks to provide richer contextual information, aiding the LLM in frame label prediction. For keyframe selection, we use the caption from the pair ($v_{t-1}$, $v_{t}$) for $v_{t}$ to maintain caption sequence coherence.
@@ 2.3 Results
Figure 19: Caption matching (left) and progress detection (right) evaluation results on HowToChange, with different VLM/LLM as evaluators.
Justification of Automatic Evaluation
Due to the lack of existing datasets with frame-wise ground truth captions, direct reference-based evaluation is infeasible. Therefore, we propose two automatic evaluation tasks, progression detection and caption matching, to assess frame-wise caption quality. To validate the reliability of these two metrics, we conduct experiments using different LLM/VLMs as evaluators for the two metrics (we pick the most widely adopted ones—Gemini and GPT for VLMs, Llama and GPT for LLMs). Figure 19 demonstrates consistent trends across these different evaluators, confirming the robustness of our evaluation methodology.